OpenStudy (anonymous):

Check all of the points that are solutions to the system of inequalities. y>4x + 3 - x>1 A.(1, 11) B.(1, 19) C.(2, 11) D.(4, 11) E.(4, 19) F.(1, -1)

OpenStudy (anonymous):

only E.

OpenStudy (anonymous):

looks like you can nix F

OpenStudy (anonymous):

why not D?

OpenStudy (anonymous):

only A, and B

OpenStudy (gw2011):

I assume the problem is: y is greater than or equal to 4x+3 because of the - below; and x>1. If this is correct, then A, B and F can be eliminated because x in each case is equal to 1 and not greater than 1. This leaves C, D and E: C) y>=(4)(2)+3 11>=8+3 11>=11 This is a true statement D) y>=(4)(4)+3 11>=16+3 11>=19 This is not a true statement E) y>=(4)(4)+3 19>=16+3 19>=19 This is a true statement Therefore, C and E would be correct if my assumption of the problem is correct